OpenStudy (anonymous):

How do you go to from the first form to the second y=4x/(x+3) at (-5,10) [4(x+h)/(x+h)+3]-[4x/(x+3)] all over h to this 12h/(x+h+3)(x+3)/h

OpenStudy (anonymous):

lots of algebra

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

it is \[f(x)=\frac{4x}{x+3}\] right?

OpenStudy (anonymous):

i tried adding just H I tried factoring out 1/h

OpenStudy (anonymous):

yes at (-5,10)

OpenStudy (anonymous):

oh you want it at a point, not in general so you want \[\frac{f(x)-10}{x-5}\]

OpenStudy (anonymous):

make that \[\frac{f(x)-10}{x+5}\]

OpenStudy (anonymous):

want to find slope at the point and equation for tan line

OpenStudy (anonymous):

f(x+h)-f(x)/h

OpenStudy (anonymous):

ok so you either want this \[\frac{f(x+h)-f(x)}{h}\] with no particular point, you can plug in the numbers later, or you want the specific one \[\frac{f(x)-10}{x+5}\]

OpenStudy (anonymous):

the second will be easier, the first is more general you can plug in \(-5\) once you find it

OpenStudy (anonymous):

I just want to know how they got 12h on top

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that =30?

OpenStudy (anonymous):

lets do it in general takes more algebra is all first of all, lets ignore the \(h\) in the denominator, and once we compute \[f(x+h)-f(x)\]we can divide by \(h\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f(x+h)=\frac{4(x+h)}{x+h+3}-\frac{4x}{x+3}\] is a start now we have to subtract

OpenStudy (anonymous):

\[\frac{4(x+h)(x+3)-4x(x+h+3)}{(x+h+3)(x+3)}\] is step 1 leave the denominator in factored form

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

all you did was multiply x+h+3 to nominator of second fraction?

OpenStudy (anonymous):

no sorry cross multiply

OpenStudy (anonymous):

i did this \[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\]

OpenStudy (anonymous):

to have common denominator

OpenStudy (anonymous):

right, that is the only way to add or subtract fractions

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now just plug -5 for x?

OpenStudy (anonymous):

now we multiply all that mess out in the top\[\frac{4(x+h)(x+3)-4x(x+h+3)}{(x+h+3)(x+3)}\] \[\frac{4x^2+4xh+12x+12h-4x^2-4xh-12x}{(x+h+3)(x+3)}\]

OpenStudy (anonymous):

which is 12

OpenStudy (anonymous):

actually it is \(12h\) but when you divide by \(h\) you are left only with the \(12\) up top

OpenStudy (anonymous):

so slope at -5 is 3

OpenStudy (anonymous):

12/(x+3)^2

OpenStudy (anonymous):

that leaves\[\frac{12}{(x+3)^2}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

at \(x=-5\) you get \(3\) it would have been lots less algebra to do it the other way

OpenStudy (anonymous):

and the line is

OpenStudy (anonymous):

you got the slope \(m=3\) and the point \((-5,10)\) point slope formula gets it

OpenStudy (anonymous):

just learning whats in book, not very good at math really cant even remember the easy stuff

OpenStudy (anonymous):

y=3x+25

OpenStudy (anonymous):

guess what? it is all algebra lots of lots of algebra

OpenStudy (anonymous):

but in a week you will be able to find the derivative of \(\frac{4x}{x+3}\) much quicker

OpenStudy (anonymous):

@satellite73 comp dropped out thanks for help. I hope so.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so that other way works the same?

OpenStudy (anonymous):

if you do it the other way you just get a number, namely 3 it is easier, but doing it this way you get a formula that you can use for any number, not just \(-5\) you will notice that we did not use \((-5,10)\) anywhere in the computation

OpenStudy (anonymous):

which would be -10 -10/x-5

OpenStudy (anonymous):

it would be \[\lim_{x\to -5}\frac{f(x)-10}{x+5}\]

OpenStudy (anonymous):

or \[\lim_{x\to -5}\frac{\frac{4x}{x+3}-10}{x+5}\]

OpenStudy (anonymous):

except for the tan line

OpenStudy (anonymous):

which would be 0/0

OpenStudy (anonymous):

right for that we used it, but once you have a general formula you can find the slope for any \(x\) so you can find the tangent line at any point

OpenStudy (anonymous):

so you still have to do the algebra

OpenStudy (anonymous):

yes, of course once you get \(\frac{0}{0}\) it just means you have more work to do

OpenStudy (anonymous):

so yo ucan use the other form for any point and sub y for f(x)

OpenStudy (anonymous):

*can't

OpenStudy (anonymous):

\[\lim_{x\to -5}\frac{\frac{4x}{x+3}-10}{x+5}\] \[=\lim_{x\to -5}\frac{{4x-10(x+3)}}{(x+5)(x+3)}\]\[=\frac{-6x-30}{(x+3)(x+5)}\] \[=\frac{-6(x+5)}{(x+3)(x+5)}=\frac{-6}{x+3}\] put in \(x=-5\) get \(3\)

OpenStudy (anonymous):

nice